### v1

# DianJin-OCR-R1: Enhancing OCR Capabilities via a Reasoning-and-Tool Interleaved Vision-Language Model

Abstract
Recent advances in large vision-language models (LVLMs) have enabled a new paradigm of end-to-end document image parsing, excelling in Optical Character Recognition (OCR) tasks such as text, table, and formula recognition. However, generative LVLMs, similarly to large language models (LLMs), are prone to hallucinations——generating words that do not exist in input images. Furthermore, LVLMs are designed for general purposes and tend to be less effective on OCR tasks compared to expert models that are trained on domain-specific datasets. In this paper, we propose DianJin-OCR-R1, a reasoning-enhanced framework designed to address these limitations through training reasoning-and-tool interleaved VLMs. Given a recognition instruction, our DianJin-OCR-R1 model first recognizes the content in the input image by its own OCR capabilities, and then calls other tools (i.e., other expert models) to obtain their results as references, finally ”looks again” the image and rethinks about the reasoning process to provide the final recognized content. Since architectures of expert models are tailored for specific OCR tasks, which makes them less prone to hallucinations, their results can help VLMs mitigate hallucinations. Additionally, expert models are typically smaller in scale and easy to iterate, enabling performance improvements for VLMs at a lower cost. We evaluate our model on ReST (Yu et al., 2023) and OmniDocBench (Ouyang et al., 2025), and experimental results show that our DianJin-OCR-R1 models consistently outperform their non-reasoning counterparts and expert OCR models, which proves the effectiveness of our method.
1 Introduction
Extracting structured content from document images containing intertwined elements such as plain texts, tables, and formulas is crucial for advancing AI-assisted document understanding, and it is also one of the most critical steps in RAG (Retrieval-Augmented Generation). Traditional expert OCR models like PP-Structure (Li et al., 2022a; Cui et al., 2025) train expert models for different tasks separately, such as SLANet (Li et al., 2022a) for table recognition and PP-FormulaNet (Liu et al., 2025) for formula recognition. Compared to PP-Structure with complex procedures, large vision-language models (LVLMs) like Qwen2.5-VL (Bai et al., 2025) can process various document elements in an unified, elegant, and end-to-end paradigm. In particular, expert VLMs such as GOT (Wei et al., 2024) are trained on various OCR tasks to directly generate structured content from document images.
Despite these advancements, we observe that current general and expert VLMs are prone to generate plausible-sounding answers that are not always grounded in the input image, a phenomenon usually referred to as hallucination. In OCR tasks, since they have learned the language patterns from the training datasets used in LLMs instead of carefully recognizing the actual content in the images, they often output content that does not exist in the images. In contrast,although expert OCR models can only accomplish one specific task, they rarely have hallucination issues because their architectures are tailored for specific OCR tasks. At worst, they may fail to recognize one or a few characters. Table 1 shows an example of hallucination. Moreover, LVLMs mainly focus on visual reasoning performance and their capabilities in perception are not as strong, while expert OCR models have stronger fine-grained perception capabilities that are beneficial for recognizing various texts.
[Uncaptioned image] Prompt 请识别图中印章的内容,直接输出内容即可。
Qwen2.5-VL-7B-Instruct 北川羌族自治县农村信用合作联社永安分社
$P$-StructureV3 北川族自治县永安工农 第九村民小组
Ground-truth 北川羌族自治县永安镇工农村 第九村民小组
Table 1: An example of hallucination
To synergize the advantages of both approaches while addressing their limitations, we present Dianjin-OCR-R1, a novel reasoning-enhanced framework following an reasoning-and-tool interleaved paradigm. Inspire by reasoning models that are trained to articulate their problem-solving process through a Chain-of-Thought (CoT) manner, our Dianjin-OCR-R1 model first recognizes the content in the input image through its own OCR capabilities, then calls other expert models or tools and uses their results as references, finally ’looks again’ the image and rethinks about the reasoning process to provide the final recognized content. The results of expert models can help VLMs reflect their own results, enabling it as less hallucinating as possible. Additionally, expert models are typically smaller in scale and easy to iterated, which allows VLMs to improve their performance at a lower cost.
We train our DianJin-OCR-R1 model based on Qwen2.5-VL-7B-Instruct (Bai et al., 2025) by Supervised Fine-Tuning (SFT) and Reinforcement Fine-Tuning (RFT). We apply Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a reinforcement learning algorithm that introduces two reward signals: a format reward to encourage structured outputs and an accuracy reward to promote answer correctness. These mechanisms guide the model to produce coherent, verifiable reasoning chains and reliable answers. Our contributions in this work are two-fold:
  • We propose a novel reasoning-and-tool interleaved framework, namely Dianjin-OCR-R1, balancing the reasoning ability of LVLMs and the OCR ability of export models. Experimental results on benchmarks including ReST (Yu et al., 2023) and OmniDocBench (Ouyang et al., 2025) show that our model consistently outperforms their non-reasoning counterparts and expert models, which proves the effectiveness of our method.
  • Our framework only requires the iteration of export OCR models (typically with parameters less than LVLMs), without need for the post-training of LVLMs, to comprehensively enhance the OCR performance, which allows us to better utilize resources.
2 Related Work
2.1 General VLMs
General VLMs such as GPT-4V (Yang et al., 2023b), Claude series [^1] https://www.anthropic.com/news/claude-3-5-sonnet, Gemini series (Team et al., 2024), Qwen-VL series (Bai et al., 2023; Wang et al., 2024c; Bai et al., 2025), MiniCPM series (Yao et al., 2024), Intern-VL series (Chen et al., 2024b; a; Zhu et al., 2025), and Seed-VL series (Guo et al., 2025b) have demonstrated encouraging results in document parsing and understanding, and do not require training for specific tasks. However, general LVLMs focus on visual reasoning performance, e.g., VQA, and their capabilities in perception are not as strong. Given this limitation, some studies argue that LVLMs are not well-suited for OCR tasks.
2.2 Expert VLMs
Expert VLMs such as GOT (Wei et al., 2024), Monkey series (Liu et al., 2024; Li et al., 2025), mPLUG-DocOwl-series (Ye et al., ; Ye et al., 2023), olmOCR (Poznanski et al., 2025), Ocean-OCR (Chen et al., 2025), and Mistral-OCR [^2] https://mistral.ai/fr/news/mistral-ocr are specifically designed and trained for document parsing and understanding tasks. They used an innovative unified model that processes various document elements. However, they still suffer from hallucinations and have high iteration costs.
2.3 Expert OCR Models
Researchers usually trained different expert OCR models for different OCR tasks separately. PP-OCR series (Du et al., 2020; 2021; Li et al., 2022b; Liu et al., 2025) focus on multi-scenario and multi-lingual text detection and text recognition tasks. TableMaster (Cao & Liu, 2025), SLANet (Li et al., 2022a), and UniTable (Peng et al., 2024) are proposed to recognize the structure and content of tables. LaTeX-OCR [^3] https://github.com/lukas-blecher/LaTeX-OCR, PP-FormulaNet (Liu et al., 2025), and UniMERNet (Wang et al., 2024a) are designed for formula recognition task. However, one inconvenient aspect is that researchers often have to switch between different models to meet various recognition requirements.
2.4 Reasoning Models
Recent advances in reasoning models, optimized through reinforcement learning, highlight the potential for foundation models to enhance inference-time scaling and achieve higher levels of intelligence, e.g., OpenAI o3444https://openai.com/index/introducing-o3-and-o4-mini, DeepSeek-R1 (Guo et al., 2025a). Following the success of text-only reasoning models, LVLM reasoning research has emerged, focusing on both effective multi-modal chain-of-thought structures and high-quality training data construction methods. However, the majority of them are developed for VQA tasks, with few intended for OCR tasks.
3 Method
In this section, we provide a comprehensive overview of DianJin-OCR-R1, delving into the details of its reasoning paradigm, data construction, and model training.
3.1 Reasoning Paradigm
Our proposed DianJin-OCR-R1, focusing on recognition tasks, fosters a straightforward reasoning paradigm in LVLMs, enabling a more transparent and verifiable reasoning process. Given an image and a textual question, DianJin-OCR-R1 first enables the model to recognize relevant content in the input image using its own OCR capability. Then, it calls on other expert models or tools, using their results as references or supplementary information. Next, it ”looks again” the image, comprehensively analyzes both its own results and those from other models, reflects on whether it made any mistakes or missed anything during recognition. Finally, the model provides the final recognized content. Figure 1 shows the visualization of the reason pipeline of our proposed method.
Figure 1: The visualization of the reason pipeline of our proposed method.
3.2 Data Construction
To prove the effectiveness of our method, we conduct experiments on three OCR tasks, i.e., seal, table, and formula recognition tasks. We denote our original seal, table, and formula recognition datasets as $D_{sr}/D_{tr}/D_{fr}=(i_{i},q_{i},a_{i})$, where $i_{i}$ denotes the i-th seal, table, or formula image, $q_{i}$ refers to the instruction prompting the model to recognize, and $a_{i}$ corresponds to the seal, table (in HTML format), or formula (in LaTeX format) content.
We leverage the advanced capabilities of powerful LVLMs, such as Qwen-VL-Max [^5] https://help.aliyun.com/zh/model-studio/vision, through a well-designed prompt and an example to generate an interleaved reasoning-and-tool chain $r_{i}$ and an output $o_{i}$. The reasoning chain $r_{i}$ consists of three parts, content recognized by model itself enclosed within <think></think> tags, multiple tool responses enclosed within <tool></tool> tags, reflects enclosed within <rethink></rethink> tags. The generated output is enclosed within <answer></answer> tags. If the output $o_{i}$ matches the ground-truth $a_{i}$, we retain the instance $(i_{i},q_{i},r_{i},o_{i})$ as a valid reasoning sample, with which we ultimately construct three reasoning datasets $R_{sr}/R_{tr}/R_{fr}=(i_{i},q_{i},r_{i},o_{i})$. Figure 2 shows the pipeline for constructing reasoning data.
Considering the differences between three tasks, we select data from different sources and use different tools when constructing reasoning chains.
Seal Recognition.
We select data from ReST (Yu et al., 2023), which is from ICDAR 2023 competition on reading the seal title, including a train set with 5,000 images and a test set of 5,000 images. Since the annotations of test set are not publicly available, we only used the train set to verify our method. We found that there is considerable repetition in the seal titles. To prevent the LLM part from memorizing fixed phrases in the seal titles, we isolate the data during splitting train and test sets (using community detection for coarse filtering and specific keywords for fine filtering). When generating the reasoning chain, we use the results of PP-StructureV3 and Qwen-VL-OCR  [^6] https://help.aliyun.com/zh/model-studio/qwen-vl-ocr as references. If the output $o$ and the answer $a$ match exactly in terms of characters, we retain the instance.
Table Recognition.
We select data from in-house datasets and TabRecSet (Yang et al., 2023a) dataset that is the largest and first bi-lingual dataset for end-to-end TR, with 38.1K tables in which 20.4K are in English and 17.7K are in Chinese. When generating the reasoning chain, we use the results of PP-StructureV3 and MonkeyOCR-3B as references. If the Tree-Edit-Distance-based Similarity (TEDS) (Zhong et al., 2020) score between the output $o$ and the answer $a$ lager than the content recognized by model itself and exceeds a threshold (set to 0.98), we retain the instance as a valid reasoning sample.
Formula Recognition.
We select data from UniMER-1M (Wang et al., 2024a) dataset, which features over one million instances that represent a diverse and intricate range of mathematical expressions. Similarly, we use the results of PP-StructureV3 and MonkeyOCR-3B as references in the reasoning chain. If the 1-Normalized Edit Distance (1-NED) (Zhang et al., 2019) score between the output $o$ and the answer $a$ smaller than the content recognized by model itself and is below a threshold (set to 0.015), we retain the instance as a valid reasoning sample.
Figure 2: Overall pipeline for constructing reasoning data.
3.3 Model Training
In this section, we present the details for training VLMs to perform OCR reasoning. The training consists of two common post-training strategies: learning reasoning with Supervised Fine-Tuning (SFT) and Reinforcement Fine-Tuning (RFT).
Learning reasoning with SFT.
The reasoning datasets $R_{sr}$, $R_{tr}$, and $R_{fr}$ are utilized to fine-tune VLMs to generate a CoT followed by a final answer. Each training instance $(p,i,r,o)$ consists of a image $i$, a question $q$, a reasoning chain $r$ and a output $o$. During SFT stage, the image $i$ and the question $q$ serve as the input to the model, while the reasoning chain $r$ and the output $o$ are treated as the target output, enabling the model to learn to produce coherent reasoning steps along with the correct solution.
Learning reasoning with RFT.
The filtered datasets $D^{{}^{\prime}}_{sr}$, $D^{{}^{\prime}}_{tr}$, and $D^{{}^{\prime}}_{fr}$ (same size and input with $R_{sr}$, $R_{tr}$, and $R_{fr}$) from $D_{sr}$, $D_{tr}$, and $D_{fr}$ are utilized to reinforcement fine-tune, which aims to enhance the reasoning skills of the base model. Specifically, we adopt the GRPO algorithm for RFT stage, incorporating two reward mechanisms: a format reward to ensure the generated output adheres to the desired structure, and an accuracy reward to encourage correct answers.
  • Format reward. We incorporate a format reward to promote well-structured outputs. For a given CoT, a reward score of 1.0 is granted if the CoT contains exactly <think></think>, <tool></tool>, <rethink></rethink> and <answer></answer> tags, with no additional content outside these boundaries. If this strict formatting criterion is not met, it receives a reward of 0.0.
  • Accuracy reward. We use an accuracy reward to promote answer correctness. Specifically, for seal recognition, if the output $o$ exactly matches the answer $a$, the model receives a reward score of 1.0; otherwise, it receives a score of 0.0. For table recognition, the TEDS score between $o$ and $a$ is used as the reward score. For formula recognition, the Character Detection Matching (CDM) (Wang et al., 2024b) metric between $o$ and $a$ serves as the reward score, and an additional 0.5 will be added if the score is equal to 1.0.
4 Experimentation
4.1 Experimental Setups
Training Data
Table 2 shows the statistics of the train sets.
Datasets size $i_{token}$ $q_{token}$ $r_{token}$ $a_{token}$
$R_{sr}$ 1024 54 106 417 14
$R_{tr}$ 1024 15 369 1697 324
$R_{fr}$ 1024 17 54 769 65
Table 2: Statistics of the train sets.
Model Training
We train DianJin-OCR-R1 based on Qwen2.5-VL-7B-Instruct. Our model training consists of two strategies, conducted on a single node with 8 NVIDIA A100 GPUs. In the SFT stage, we leverage DeepSpeed’s Zero-3 for optimization, using a learning rate of $1.0\times 10^{-5}$, a sequence length of 16K, and bf16 precision over 2 epochs. We apply gradient accumulation with 4 steps to simulate a larger batch size. In the RFT stage, we perform 8 rollouts per sample, with a train batch size of 128 and a rollout batch size of 32. We use a learning rate of $1.0\times 10^{-6}$ and a sampling temperature of 1.0 over 3 epochs to balance learning dynamics.
Evaluation Datasets
We evaluated our models on three benchmarks, including the ReST test sets, the OmniDocBench table sets, and the OmniDocBench formula sets. They include multiple languages, such as Chinese, English, and a mix of both. The detailed statistics of these test sets are summarized in Table 3. For ReST testing sets, the results are evaluated by accuracy. For OmniDocBench table sets, we report the TESD, structure TEDS (STEDS), and Normalized Edit Distance. For OmniDocBench formula sets, we evaluate the results using CDM, ExpRate@CDM (Wang et al., 2024b) and Normalized Edit Distance.
Dataset Language Size
ReST Test* Chinese 432
OmniDocBench Table Chinese/English/Mixed 428
OmniDocBench Formula Chinese/English 353
Table 3: Statistics of the test sets.
4.1.1 Baselines
We compare our model against three categories of models. The first category includes General LVLMs: Qwen2.5-VL-7B-Instruct (Bai et al., 2025), and Intern3VL-8B (Zhu et al., 2025). The second category consists of Expert VLMs, including Qwen-VL-OCR, GOT (Wei et al., 2024), and MonkeyOCR-3B (Li et al., 2025). The third category includes Expert OCR Models: PP-StructureV3 (PP-OCRv4_server_seal_det + PP-OCRv4_server_rec) for seal recognition, PP-StructureV3 (PP-LCNet_x1_0_table_cls + SLANeXt_wired + SLANeXt_wireless) and RapidTable [^7] https://github.com/RapidAI/RapidTable (SLANet_plus) for table recognition, and PP-StructureV3 (PP-FormulaNet_plus-M) and UniMERNet-B (Wang et al., 2024a) for formula recognition.
4.2 Experimental Results
Table 4 compare the performance of our DinJin-OCR-R1 model with other models. The results reveal the following key points:
  • On the three test sets, our DianJin-OCR-R1 models significantly outperform the base model (Qwen2.5-VL-7B-Instruct). For instance, our DianJin-OCR-R1 model trained by SFT improves accuracy from 0.527 to 0.722 on seal recognition task, TEDS from 0.858 to 0.895 on table recognition task, and CDM from 0.915 to 0.977 on formula recognition task.
  • Our DianJin-OCR-R1 models outperform the expert OCR models using in the reasoning chains, which indicates that they can integrate the advantages of various models.
  • We achieve best results with the model using only thousands of training samples. Under the same data conditions, the models trained by RFT are generally superior to those trained by SFT, indicating that RFT is more advantageous.
Model Seal Table Formula
[…]
Expert OCR Models
$P$-StructureV3 0.600 0.818 0.918 0.121 0.966 0.730 0.198
RapidTable - 0.832 0.902 0.120 - - -
UniMERNet-B - - - - 0.932 0.645 0.290
General VLMs
Qwen2.5-VL-7B-Instruct 0.527 0.853 0.904 0.107 0.915 0.642 0.243
InternVL3-8B 0.475 0.812 0.880 0.139 0.946 0.668 0.273
Expert VLMs
GOT - 0.787 0.885 0.219 0.898 0.622 0.287
MonkeyOCR-3B - 0.845 0.895 0.182 0.973 0.767 0.196
Qwen-VL-OCR 0.472 - - - - - -
Dianjin-OCR-R1
Dianjin-OCR-R1(by SFT) 0.722 0.895 0.935 0.080 0.977 0.770 0.188
Dianjin-OCR-R1(by RFT) 0.766 0.901 0.939 0.072 0.976 0.787 0.179
Table 4: Performance comparison across different test sets. Scores in bold and underlined indicate the best and second-best results, respectively.
4.3 Ablation Study
To assess the contribution of each component in our DianJin-OCR-R1 framework, we conduct comprehensive ablation experiments. Table 5 summarizes our findings.
  • Baseline vs Baseline with Tools. Directly supplying the outputs of two tools as references to the baseline Qwen2.5-VL-7B-Instruct significantly boosts its performance (e.g., increasing accuracy from 0.551 to 0.708 on seal recognition tasks), which demonstrates that leveraging expert models can effectively enhance the capabilities of VLMs on OCR tasks.
  • Fine-tuning directly vs Fine-tuning with Reasoning and Tools. We conducted direct SFT and RFT on the base model, using only the image $i$ and question $a$ as inputs and the answer $a$ as output, omitting the reasoning chain $r$. The results indicate that models incorporating reasoning chains and tools consistently achieve superior performance, highlighting the successful integration of reasoning and tools.
  • Baseline with tools vs Fine-tuning with Reasoning and Tools. Comparing to the baseline with two tools, models trained by SFT with reasoning and tools perform better, which demonstrates the effectiveness of distillation method.
Model Seal Table Formula
[…]
Baseline 0.527 0.853 0.904 0.107 0.915 0.642 0.243
Baseline w tools 0.560 0.865 0.908 0.107 0.958 0.713 0.241
SFT 0.643 0.863 0.910 0.098 0.938 0.599 0.196
SFT w tools 0.722 0.895 0.935 0.080 0.977 0.770 0.188
RFT 0.634 0.897 0.937 0.078 0.974 0.727 0.204
RFT w tools 0.766 0.901 0.939 0.072 0.976 0.787 0.179
Table 5: Ablation study
Furthermore, we also replace the tools to conduct more in-depth exploration. For formula recognition task, we replace MonkeyOCR-3B and PP-FormulaNet_plus-M with GOT and PP-FormulaNet_plus-S. As shown in Table 6, the performance using MonkeyOCR-3B and PP-FormulaNet_plus-M are better than using GOT and PP-FormulaNet_plus-S, which demonstrates that optimizing a expert OCR model alone whose costs are lower can improve the OCR capabilities of our model.
[…]
Tool1: GOT 0.898 0.622 0.287
Tool2: PP-FormulaNet_plus-S 0.852 0.523 0.299
DianJin-OCR-R1 (w tools) 0.969 0.722 0.227
Tool1: MonkeyOCR-3B 0.973 0.767 0.196
Tool2: PP-FormulaNet_plus-M 0.966 0.730 0.197
DianJin-OCR-R1 (w tools) 0.979 0.770 0.189
Table 6: Performance comparison of models using different tools.
5 Conclusion
We have presented DianJin-OCR-R1, a reasoning-augmented framework for large vision-language models on OCR tasks. This framework fosters a straightforward reasoning paradigm in VLMs, enabling a more transparent and verifiable reasoning process on complex document OCR tasks. Through extensive experiments on diverse benchmarks, we demonstrate that reasoning-aware training significantly improves both performance.
Acknowledgements
We extend our sincere thanks to other members of the Qwen DianJin Team for their significant contributions, whose hard work and dedication were crucial to the success of this project.
Appendix A Prompts used in this paper
As shown in Figure 3, Figure 4 and Figure 5, we use a multi-turn conversation approach to generate data. Specifically, in the first turn, we let the LVLM recognize the contents of the image by itself. In the second turn, we provide results from other tools as references, allowing the model to analyze and give the final answer.
Prompt Template
## Turn 1
你是一名专业的OCR专家,你的任务是识别图片中的印章抬头。印章抬头位于印章上方,通常用于标明印章的归属、用途或机构名称,是印章的重要组成部分,帮助识别印章的合法性和适用范围。请将你的识别结果放在<recognition></recognition>之间。
## Turn 2
以下是其它工具对该印章的识别内容:
<tool>
    "ocr_tool_1": "{tool1}",
    "ocr_tool_2": "{tool2}"
</tool>
请比较你初步识别的结果和其它工具识别内容之间的**全部**不同之处。当存在差异时,你需要重新查看图片,并分析你的初步识别结果是否错误或者是否有遗漏。(你跟其他 OCR 工具均有可能犯错。请务必再次仔细查看图片)
将你的对比与分析过程放在<rethink></rethink>之间,并将最终识别的**印章抬头**内容放在<answer></answer>之间。
Figure 3: Prompt used to construct reasoning data for seal recognition.
Prompt Template
## Turn 1
请将图中的表格以 HTML 格式输出。
## Turn 2
以下是其它工具识别的 HTML 格式的内容:
<tool>
## 工具1:
{tool1}
## 工具2:
{tool2}
</tool>
请重新仔细查看图片,比较之前识别的结果和其它工具结果,分析表格结构(如单元格是否和图中一致、合并单元格 colspan 和 rowspan 是否正确等)和表格内容(如单元格中的内容是否识别错误或有遗漏)。如果有合并单元格,请以 colspan="x"/rowspan="x" 表示。
将分析过程放在<rethink></rethink>之间,并将最终识别的 HTML 放在<answer></answer>之间。
## 输出格式:
<rethink>
</rethink>
<answer>
HTML
</answer>
Figure 4: Prompt used to construct reasoning data for table recognition.
Prompt Template
## Turn 1
请用 LaTeX 格式写出图中公式的表达式。
## Turn 2
以下是其它工具识别的 LaTeX 格式的内容:
<tool>
## 工具1结果:
{tool1}
## 工具2结果:
{tool2}
</tool>
请重新仔细查看图片,比较之前识别的结果和其它工具结果之间的不同之处,分析是否识别错误或有遗漏。
将分析过程放在 <rethink></rethink> 之间,并将最终识别的 LaTeX 放在 <answer></answer> 之间。
## 输出格式:
<rethink>
</rethink>
<answer>
LaTeX
</answer>
Figure 5: Prompt used to construct reasoning data for formula recognition.

### v2

@@
 Since architectures of expert models are tailored for specific OCR tasks, which makes them less prone to hallucinations, their results can help VLMs mitigate hallucinations.
-Additionally, expert models are typically smaller in scale and easy to iterate, enabling performance improvements for VLMs at a lower cost.
-We evaluate our model on ReST (Yu et al., 2023) and OmniDocBench (Ouyang et al., 2025), and experimental results show that our DianJin-OCR-R1 models consistently outperform their non-reasoning counterparts and expert OCR models, which proves the effectiveness of our method.
+We evaluate our model on ReST and OmniDocBench, and experimental results show that our DianJin-OCR-R1 models consistently outperform their non-reasoning counterparts and expert OCR models, which proves the effectiveness of our method.
+Additionally, the results indicate that enhancing expert models, which are typically small and easy to iterate, enable performance improvements for VLMs.
 1 Introduction
@@
 4.1 Experimental Setups
-Training Data
+Training Datasets
 Table 2 shows the statistics of the train sets.
-Datasets size $i_{token}$ $q_{token}$ $r_{token}$ $a_{token}$
+Datasets size $q_{token}$ $i_{token}$ $r_{token}$ $a_{token}$
 $R_{sr}$ 1024 54 106 417 14
@@
 Table 3: Statistics of the test sets.
-4.1.1 Baselines
+Baselines
 We compare our model against three categories of models.
@@
 Table 4 compare the performance of our DinJin-OCR-R1 model with other models. The results reveal the following key points:
-  • On the three test sets, our DianJin-OCR-R1 models significantly outperform the base model (Qwen2.5-VL-7B-Instruct).
+  • Our DianJin-OCR-R1 models significantly outperform the base model (Qwen2.5-VL-7B-Instruct) on the three test sets.
 For instance, our DianJin-OCR-R1 model trained by SFT improves accuracy from 0.527 to 0.722 on seal recognition task, TEDS from 0.858 to 0.895 on table recognition task, and CDM from 0.915 to 0.977 on formula recognition task.
   • Our DianJin-OCR-R1 models outperform the expert OCR models using in the reasoning chains, which indicates that they can integrate the advantages of various models.
-  • We achieve best results with the model using only thousands of training samples.
+  • Our DianJin-OCR-R1 models achieve best results with the model using only thousands of training samples.
 Under the same data conditions, the models trained by RFT are generally superior to those trained by SFT, indicating that RFT is more advantageous.
@@
   • Baseline vs Baseline with Tools.
-Directly supplying the outputs of two tools as references to the baseline Qwen2.5-VL-7B-Instruct significantly boosts its performance (e.g., increasing accuracy from 0.551 to 0.708 on seal recognition tasks), which demonstrates that leveraging expert models can effectively enhance the capabilities of VLMs on OCR tasks.
+Directly supplying the outputs of two tools as references to the baseline Qwen2.5-VL-7B-Instruct significantly boosts its performance (e.g., increasing accuracy from 0.527 to 0.560 on the seal recognition task), which demonstrates that leveraging expert models can effectively enhance the capabilities of VLMs on OCR tasks.
   • Fine-tuning directly vs Fine-tuning with Reasoning and Tools.
-We conducted direct SFT and RFT on the base model, using only the image $i$ and question $a$ as inputs and the answer $a$ as output, omitting the reasoning chain $r$.
-The results indicate that models incorporating reasoning chains and tools consistently achieve superior performance, highlighting the successful integration of reasoning and tools.
+Comparing to models trained by SFT and RFT directly (only using the answer $a$ as output and omitting the reasoning chain $r$), our models incorporating reasoning chains and tools consistently achieve superior performance, which highlights the successful integration of reasoning and tools.
   • Baseline with tools vs Fine-tuning with Reasoning and Tools.
-Comparing to the baseline with two tools, models trained by SFT with reasoning and tools perform better, which demonstrates the effectiveness of distillation method.
+Comparing to the baseline with two tools, our models trained by SFT with reasoning and tools perform better, which demonstrates the effectiveness of distillation method.
 Model Seal Table Formula
@@
 RFT w tools 0.766 0.901 0.939 0.072 0.976 0.787 0.179
-Table 5: Ablation study
+Table 5: Performance comparison of models using different training strategies.
 Furthermore, we also replace the tools to conduct more in-depth exploration.
 For formula recognition task, we replace MonkeyOCR-3B and PP-FormulaNet_plus-M with GOT and PP-FormulaNet_plus-S.
-As shown in Table 6, the performance using MonkeyOCR-3B and PP-FormulaNet_plus-M are better than using GOT and PP-FormulaNet_plus-S, which demonstrates that optimizing a expert OCR model alone whose costs are lower can improve the OCR capabilities of our model.
+As shown in Table 6, the performance using MonkeyOCR-3B and PP-FormulaNet_plus-M are better than using GOT and PP-FormulaNet_plus-S, which demonstrates that optimizing a expert OCR model alone whose costs are lower can improve the OCR capabilities of our models.
 […]
@@
 Tool1: MonkeyOCR-3B 0.973 0.767 0.196
-Tool2: PP-FormulaNet_plus-M 0.966 0.730 0.197
-DianJin-OCR-R1 (w tools) 0.979 0.770 0.189
+Tool2: PP-FormulaNet_plus-M 0.966 0.730 0.198
+DianJin-OCR-R1 (w tools) 0.977 0.770 0.188
 Table 6: Performance comparison of models using different tools.
@@
 This framework fosters a straightforward reasoning paradigm in VLMs, enabling a more transparent and verifiable reasoning process on complex document OCR tasks.
-Through extensive experiments on diverse benchmarks, we demonstrate that reasoning-aware training significantly improves both performance.
+Through extensive experiments on diverse benchmarks, we demonstrate that reasoning-aware training significantly improves performance.
 Acknowledgements